In NYC, the A train takes 28 minutes to return to the station. The B train takes 42 minutes to return to the station. If both trains leave the station at the same time, how long will take for both trains to return to station at the same time?
If both trains leave the station at 9:00 am, what time will they both leave the station together again?

Answer by ikleyn(52865)

.
Least Common Multiple (LCM) of the numbers  28 = 14*2  and  42 = 14*3  is  14*6 = 84 minutes.


Answer.  It will happen every 84 minutes.